• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EST 10:08
CET 16:08
KST 00:08
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Finals Preview: Sweet Victory!12[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5
Community News
Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win0Weekly Cups (October 19-25): Clem triples first in 5.0.175Evil Geniuses - shutting down23ISSL (IntoTheTV x SOOP SC2 League): Premier7Weekly Cups (October 12-18): 5.0.17 era begins1
StarCraft 2
General
No more SC2 streamers on team liquid? [D] Hacker peak in sc2 ? Liquibets sc2 season 30 Evil Geniuses - shutting down Weekly Cups (Oct 26 - Nov 1): MaxPax, herO, Solar win
Tourneys
Sparkling Tuna Cup - Weekly Open Tournament WardiTV Mondays ISSL (IntoTheTV x SOOP SC2 League): Premier Stellar Fest TWO the Moon (Dec 16-20) RSL goes to London! 2026 Offline Finals Nov 21-22
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
Starcraft 2 Advance [D] the control group Simple Questions/Answers [G] How to playtest arcade maps
External Content
Mutation # 550 Graveyard Shift The PondCast: SC2 News & Results Mutation # 549 Endless Sparkles Mutation # 548 Shining Bright
Brood War
General
BGH Auto Balance -> http://bghmmr.eu/ Sagi.gg Launcher Released Top Players of October 2026 (EBPD Rankings) BW General Discussion Rollback Netcode in Brood War
Tourneys
[Megathread] Daily Proleagues KCM Race Survival 2026 Season 4 [Mixed-Gender CK] Rush+4 vs Best+4 [BSL23] RO24 Group F - Sunday 21:00 CET
Strategy
50,000-Word Terran Builds Memo 2026-10-22 Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers
Other Games
General Games
Stormgate/Frost Giant Megathread Nintendo Switch Thread Deus Ex 3!? Mobile Legends: Bang Bang Armies of Exigo - YesYes?
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
US Politics Mega-thread Russo-Ukrainian War Thread Canadian Politics Mega-thread Things Aren’t Peaceful in Palestine The Chess Thread
Fan Clubs
Fan Club Resource Thread Reynor Fan Club The Flash Fan Club
Media & Entertainment
Anime Discussion Thread Movie Discussion! [Manga] One Piece
Sports
Football (Soccer) Thread TeamLiquid Health and Fitness Initiative For 2023 MLB/Baseball 2023
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
Why I built Repalyzer (1.1.…
BlueStar
Innovate or Game Over?
TrAiDoS
ShieldBattery on Mac
Kraekkling
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
Can Bots Beat Pros?? Starcr…
namkraft
Customize Sidebar...

Website Feedback

Closed Threads



Active: 9786 users

Probability Question

Blogs > SmoKing2012
Post a Reply
1 2 Next All
SmoKing2012
Profile Blog Joined January 2007
United States385 Posts
May 03 2009 23:34 GMT
#1
For the savvy..

For the final exam in one of my classes, the teacher gave us a list of 5 potential essay questions. She's going to pick 3 of them to put on the test, and we have to answer one of them. If I pick 2 essay questions to study, what are the chances that at least one of the questions I studied shows up on the test?

*
How do you like them apples, ho-bag? And how do you like those very same apples, Eggars!
EGMachine
Profile Blog Joined February 2006
United States1643 Posts
May 03 2009 23:39 GMT
#2
figure it out yourself, its not our test.
I'm like, the coolest
DeathSpank
Profile Blog Joined February 2009
United States1029 Posts
May 03 2009 23:44 GMT
#3
study four of them. /thread
yes.
Jonoman92
Profile Blog Joined September 2006
United States9111 Posts
Last Edited: 2009-05-03 23:57:31
May 03 2009 23:44 GMT
#4
Oh now that I see Myrmidon's post I think that's right, 90%

On May 04 2009 08:44 DeathSpank wrote:
study four of them. /thread


Well the most he would need to study in order to definitely get one he studied for would be three...
SmoKing2012
Profile Blog Joined January 2007
United States385 Posts
May 03 2009 23:47 GMT
#5
On May 04 2009 08:39 Machine[USA] wrote:
figure it out yourself, its not our test.


Thank you for your completely useless contribution.

I thought ppl here liked math problems..
How do you like them apples, ho-bag? And how do you like those very same apples, Eggars!
Myrmidon
Profile Blog Joined December 2004
United States9452 Posts
May 03 2009 23:50 GMT
#6
I think it's 90%.

The probability you get 1 or 2 questions that you studied is 1 minus the probability that you get 0 questions you studied. The probability you get unlucky and get 0 questions is the probability that...

Imagine she is choosing the 3 questions to put on your exam in order:
First question has 3/5 chance of being bad for you.
Given that the first question you didn't study, the second question has 2/4 chance of being bad.
Given that first two questions you didn't study, the third question has 1/3 chance of being bad.

So the probability you get no questions you studied is (3/5)(2/4)(1/3) = 0.1

Probability you get at least one question you studied is 1 - 0.1 = 0.9 = 90%.
SmoKing2012
Profile Blog Joined January 2007
United States385 Posts
May 03 2009 23:54 GMT
#7
On May 04 2009 08:50 Myrmidon wrote:
I think it's 90%.

The probability you get 1 or 2 questions that you studied is 1 minus the probability that you get 0 questions you studied. The probability you get unlucky and get 0 questions is the probability that...

Imagine she is choosing the 3 questions to put on your exam in order:
First question has 3/5 chance of being bad for you.
Given that the first question you didn't study, the second question has 2/4 chance of being bad.
Given that first two questions you didn't study, the third question has 1/3 chance of being bad.

So the probability you get no questions you studied is (3/5)(2/4)(1/3) = 0.1

Probability you get at least one question you studied is 1 - 0.1 = 0.9 = 90%.


Looks good at a glance, thanks.
How do you like them apples, ho-bag? And how do you like those very same apples, Eggars!
n.DieJokes
Profile Blog Joined November 2008
United States3443 Posts
May 03 2009 23:55 GMT
#8
cool, binomial probability question. So we do, 1-binomialcdf (n,p,k-1) or (5,.4,0) so .92224. If I screwed up anywhere someone call me on it, it's been a while.
MyLove + Your Love= Supa Love
Myrmidon
Profile Blog Joined December 2004
United States9452 Posts
Last Edited: 2009-05-03 23:59:17
May 03 2009 23:58 GMT
#9
On May 04 2009 08:54 SmoKing2012 wrote:
Show nested quote +
On May 04 2009 08:50 Myrmidon wrote:
I think it's 90%.

The probability you get 1 or 2 questions that you studied is 1 minus the probability that you get 0 questions you studied. The probability you get unlucky and get 0 questions is the probability that...

Imagine she is choosing the 3 questions to put on your exam in order:
First question has 3/5 chance of being bad for you.
Given that the first question you didn't study, the second question has 2/4 chance of being bad.
Given that first two questions you didn't study, the third question has 1/3 chance of being bad.

So the probability you get no questions you studied is (3/5)(2/4)(1/3) = 0.1

Probability you get at least one question you studied is 1 - 0.1 = 0.9 = 90%.


Looks good at a glance, thanks.


Yeah I hope I didn't make a fool of myself, too. I think an alternative way to think of it is there being (5 choose 3) combinations of possible exam questions. (5 choose 3) = 5! / [(3)!*(5-3)!] = 120 / [6 * 2] = 10. Only 1 of the 10 combinations is bad for you: the one that contains all 3 test questions you didn't study.
Lemonwalrus
Profile Blog Joined August 2006
United States5465 Posts
Last Edited: 2009-05-04 00:50:20
May 04 2009 00:48 GMT
#10
Since there are so few possible combinations, I just solved it this way.
Possible Essay Questions: A,B,C,D,E.
Possible combinations: (Made really simple since order doesn't matter)
ABC BCD CDE
ABD BCE
ABE BDE
ACD
ACE
ADE

Therefore, assuming you studied A and B, all of the outcomes except for CDE would have one essay that you were prepared for, so you have 90% chance to encounter one. I'd suggest at least familiarizing yourself with a third one though just in case.
Crunchums
Profile Blog Joined December 2008
United States11144 Posts
May 04 2009 00:49 GMT
#11
On May 04 2009 08:34 SmoKing2012 wrote:
For the savvy..

For the final exam in one of my classes, the teacher gave us a list of 5 potential essay questions. She's going to pick 3 of them to put on the test, and we have to answer one of them. If I pick 2 essay questions to study, what are the chances that at least one of the questions I studied shows up on the test?

1 - 1/(5 choose 2) = 9/10 = 90%
brood war for life, brood war forever
Racenilatr
Profile Blog Joined August 2008
United States2756 Posts
May 04 2009 01:11 GMT
#12
On May 04 2009 08:34 SmoKing2012 wrote:
For the savvy..

For the final exam in one of my classes, the teacher gave us a list of 5 potential essay questions. She's going to pick 3 of them to put on the test, and we have to answer one of them. If I pick 2 essay questions to study, what are the chances that at least one of the questions I studied shows up on the test?



isn't it 40%? coz it's the chance of ONE question and your studying 2/5 of the potential questions right?
Pawsom
Profile Blog Joined February 2009
United States928 Posts
May 04 2009 02:21 GMT
#13
On May 04 2009 10:11 Racenilatr wrote:
Show nested quote +
On May 04 2009 08:34 SmoKing2012 wrote:
For the savvy..

For the final exam in one of my classes, the teacher gave us a list of 5 potential essay questions. She's going to pick 3 of them to put on the test, and we have to answer one of them. If I pick 2 essay questions to study, what are the chances that at least one of the questions I studied shows up on the test?



isn't it 40%? coz it's the chance of ONE question and your studying 2/5 of the potential questions right?


no
n.DieJokes
Profile Blog Joined November 2008
United States3443 Posts
May 04 2009 02:22 GMT
#14
On May 04 2009 08:55 n.DieJokes wrote:
cool, binomial probability question. So we do, 1-binomialcdf (n,p,k-1) or (5,.4,0) so .92224. If I screwed up anywhere someone call me on it, it's been a while.

I checked it, this is the answer.
MyLove + Your Love= Supa Love
foppa
Profile Blog Joined July 2008
Canada451 Posts
May 04 2009 02:36 GMT
#15
there is a 70% chance that you will have prepared for one of the essays.

3/5 chance that you have not prepared the essays
if the teacher does a random draw:
on the first draw there will be a 3/5 chance that you have not prepared that question. on the second draw there will be a 2/4 chance you have not prepared. thus the probability of getting 2 draws of "no" will be 3/5 * 2/4 = 6/20 or 3/10 or 30%
thus knowing that the probability of knowing both questions + the probability of knowing one of the questions will be equal to 70%.

I would advice you to not be a lazy piece of crap and study for 4 as there will be a 100% chance of you knowing at least 1 of the questions on the exam.

if you study for 3 the probability of getting a question you can answer will be of 90%.

make your choice wisely.

/good luck

ps: i have a math exam next week and probability is one of the units i gotta know so you can believe my post ^-^

i can take you
DeathSpank
Profile Blog Joined February 2009
United States1029 Posts
May 04 2009 02:39 GMT
#16
wait I misread I thought he had to answer all three. Since you only have to answer 1 it would be safe to study three but you could probably pull off only studying two. Either way stop reading tl and study!
yes.
Lemonwalrus
Profile Blog Joined August 2006
United States5465 Posts
May 04 2009 02:45 GMT
#17
On May 04 2009 11:36 foppa wrote:
there is a 70% chance that you will have prepared for one of the essays.

3/5 chance that you have not prepared the essays
if the teacher does a random draw:
on the first draw there will be a 3/5 chance that you have not prepared that question. on the second draw there will be a 2/4 chance you have not prepared. thus the probability of getting 2 draws of "no" will be 3/5 * 2/4 = 6/20 or 3/10 or 30%
thus knowing that the probability of knowing both questions + the probability of knowing one of the questions will be equal to 70%.

I would advice you to not be a lazy piece of crap and study for 4 as there will be a 100% chance of you knowing at least 1 of the questions on the exam.

if you study for 3 the probability of getting a question you can answer will be of 90%.

make your choice wisely.

/good luck

ps: i have a math exam next week and probability is one of the units i gotta know so you can believe my post ^-^


There are THREE draws, not two.
foppa
Profile Blog Joined July 2008
Canada451 Posts
May 04 2009 02:49 GMT
#18
On May 04 2009 08:55 n.DieJokes wrote:
cool, binomial probability question. So we do, 1-binomialcdf (n,p,k-1) or (5,.4,0) so .92224. If I screwed up anywhere someone call me on it, it's been a while.


wrong ^-^
you can only use binomialcdf when success or failure is determined randomly with the same probability of success each time the event occurs
in this case if you pick one essay whether its a yes or a no the probability will be out of 5 (2/5 success, 3/5 failure), the second draw will be out of 4 since one essay been eliminated by the first draw

i can take you
Jonoman92
Profile Blog Joined September 2006
United States9111 Posts
May 04 2009 02:55 GMT
#19
On May 04 2009 09:48 Lemonwalrus wrote:
Since there are so few possible combinations, I just solved it this way.
Possible Essay Questions: A,B,C,D,E.
Possible combinations: (Made really simple since order doesn't matter)
ABC BCD CDE
ABD BCE
ABE BDE
ACD
ACE
ADE

Therefore, assuming you studied A and B, all of the outcomes except for CDE would have one essay that you were prepared for, so you have 90% chance to encounter one. I'd suggest at least familiarizing yourself with a third one though just in case.


Once again, simplicity and common sense wins.
foppa
Profile Blog Joined July 2008
Canada451 Posts
May 04 2009 03:00 GMT
#20
On May 04 2009 11:45 Lemonwalrus wrote:
Show nested quote +
On May 04 2009 11:36 foppa wrote:
there is a 70% chance that you will have prepared for one of the essays.

3/5 chance that you have not prepared the essays
if the teacher does a random draw:
on the first draw there will be a 3/5 chance that you have not prepared that question. on the second draw there will be a 2/4 chance you have not prepared. thus the probability of getting 2 draws of "no" will be 3/5 * 2/4 = 6/20 or 3/10 or 30%
thus knowing that the probability of knowing both questions + the probability of knowing one of the questions will be equal to 70%.

I would advice you to not be a lazy piece of crap and study for 4 as there will be a 100% chance of you knowing at least 1 of the questions on the exam.

if you study for 3 the probability of getting a question you can answer will be of 90%.

make your choice wisely.

/good luck

ps: i have a math exam next week and probability is one of the units i gotta know so you can believe my post ^-^


There are THREE draws, not two.


im gonna fail xD yes you are right I calculated the outcome of the first 2 draws. why am I doing math at 11 at night is beyond me

the 3rd draw indeed gives you a probability of a 90% of having studied for at least 1 of the essays
i can take you
1 2 Next All
Please log in or register to reply.
Live Events Refresh
OSC
23:00
King of the Hill #267
EmpressLilyy65
Liquipedia
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
LamboSC2 205
SHIN 129
BRAT_OK 67
ByuN (SOOP) 28
EnDerr 4
StarCraft: Brood War
Jaedong 1727
Horang2 1224
Mini 667
Bisu 597
Shuttle 584
Light 539
Soulkey 404
BeSt 389
Stork 356
Soma 315
[ Show more ]
ggaemo 291
actioN 248
Snow 245
ZerO 199
firebathero 176
Hyuk 173
hero 165
Sharp 113
Mong 110
ProTech110
Mind 104
Hyun 93
Bonyth 91
Leta 89
Zeus 78
Pusan 54
Killer 46
Bale 38
Shinee 36
Shine 30
sorry 30
JYJ 25
HiyA 24
Hm[arnc] 23
ajuk12(nOOB) 17
Sacsri 17
Terrorterran 17
Aegong 16
Bishop 15
soO 15
zelot 15
910 13
Free 11
scan(afreeca) 11
Dota 2
qojqva1909
Dendi972
BananaSlamJamma315
monkeys_forever279
XcaliburYe155
League of Legends
Reynor110
Counter-Strike
kRYSTAL_43
Heroes of the Storm
crisheroes315
Other Games
B2W.Neo943
Liquid`RaSZi773
Lowko429
syndereN293
Fuzer 258
XaKoH 150
Mew2King118
QueenE72
ArmadaUGS62
NeuroSwarm59
Railgan26
Liquid`VortiX22
Chillindude19
ZerO(Twitch)14
Organizations
Counter-Strike
PGL275
[ Show 12 non-featured ]
StarCraft 2
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Laughngamez YouTube
• Migwel
StarCraft: Brood War
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
Dota 2
• C_a_k_e 5290
• WagamamaTV618
League of Legends
• TFBlade808
Upcoming Events
IntoTheTV X SOOP
1d 19h
Reynor vs Solar
ByuN vs SHIN
Clem vs MaxPax
herO vs Rogue
IPSL
2 days
WorsT vs nOmaD
Dewalt vs andersOn
AI Arena Tournament
2 days
IntoTheTV X SOOP
2 days
IPSL
3 days
Azhi_Dahaki vs ggnore13
Double vs WIZARD
Replay Cast
3 days
Replay Cast
4 days
Sparkling Tuna Cup
4 days
WardiTV Invitational
4 days
YoungYakov vs Lambo
SHIN vs TBD
Percival vs Shameless
Solar vs TBD
PiGosaur Cup
5 days
[ Show More ]
The PondCast
6 days
KCM Race Survival
6 days
WardiTV Invitational
6 days
Liquipedia Results

Completed

Proleague 2026-11-04
Blizzard Classic Cup 2026
HCC Season 3

Ongoing

Acropolis #5
KCM Race Survival 2026 Season 4
International ShieldBattery League
BSL Season 23
IPSL Winter 2026-27
Calamity Cup Div. A Season 8
IEM Beijing 2026
PGL Masters Bucharest 2026
Stake Ranked Episode 5
1win Private Club #2
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026

Upcoming

1xBet Invitational
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
Underdog Cup 2026 Div 2-3
PGL Bucharest 2027
BLAST Bounty Winter 2027
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.